Operation Summary for October 12, 2009 - October 14, 2009
Monday Oct. 12
The day shift began with a Operations monitoring the Tevatron (TeV) store 7258 (194.27E30) and the Recycler stash (74.9E10), with the Antiproton Source stacking (22.3E10), with the Linac, Booster, and Main Injector (MI) operating normally, with the NuMI and MiniBooNE experiments taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) experiments using the colliding beams.
During the day shift, store 7258 collided while the Recycler stashed antiprotons.
During the evening shift, Operations terminated store 7258 at 5 PM. CDF and D0 accessed their collision halls. Operations conducted TeV orbit checks and then put the TeV into shot setup at 9 PM. Operations established HEP store 7261 with an average initial luminosity (AIL) of 259E30 at 11 PM.
Tuesday Oct. 13
The midnight shift began with Operations monitoring store 7261 (202.01E30) and the stash (73.2E10), with Pbar stacking (13.4E10), and with NuMI and MiniBooNE taking beam.
During the day shift, Operations terminated store 7261 at 1 PM. TeV experts then conducted studies until 2:30 PM. CDF and D0 accessed their collision halls through to the end of the shift.
During the evening shift, Operations conducted TeV orbit checks until 6 PM and then put the TeV into shot setup. Operations established store 7264 with an AIL of 282.8E30 at 7:49 PM.
Wednesday Oct. 14
The midnight shift began with Operations monitoring store 7264 (136.67E30) and the stash 170.6E10), with Pbar stacking (17.9E10), and with NuMI and MiniBooNE taking beam.
During the midnight shift, Operations terminated store 7264 at 6:30 AM and then put the TeV into shot setup.
The Plans for Wednesday
The plans for today are to stack, stash, and store. The Run Coordinator will allow a two-hour access into Booster, MI, NuMI and MiniBooNE this morning.
Comments and Suggestions
|last modified 10/14/2009 email Fermilab|
|Security, Privacy, Legal|